# Multi-scale tests of independence powerful for detecting explicit or implicit functional relationship

### v1

@@ 4 Performance on simulated datasets
It can be observed from this figure that $\Psi_{n}^{phi}$ performs best or next to best in the absence of noise. Even in the presence of noise, power of $\Psi_{n}^{phi}$ is quite good. As the noise level increases, $\Psi_{n}^{dcor}$ takes the lead. The overall performance of $\Psi_{n}^{cor}$ is competitive except for the ‘Circle $\lambda$’ case. HHG performed well but it lagged behind our proposed tests most of the time. HSIC performed well in ‘Parabola $\lambda$’ and ‘Circle $\lambda$’ distributions, but didn’t achieve satisfactory power in other distributions. dCor had a somewhat competitive power in ‘Parabola $\lambda$’, but not in other distributions. MIC had low power overall, except for higher sample sizes in the ‘Sine $\lambda$’ distributions. Xicor also performed poorly except for ‘Sine $\lambda$’, where it performed the best. This is not unexpected from xicor, as it is mainly designed to detect whether $Y$ is a function of $X$. BET had the lowest power among all tests in most distributions.
@@ 5 Discussion and conclusion
Figure 8: The scatter plots of ‘doppler’, ‘$BEX_{2}$’, ‘Lissajous A’, ‘Lissajous B’, ‘rose curve’, ‘spiral’, ‘tilted square’ and ‘five clouds’ distributions.

### v2

@@ 4 Performance on simulated datasets
It can be observed from this figure that $\Psi_{n}^{phi}$ performs best or next to best in the absence of noise. Even in the presence of noise, the power of $\Psi_{n}^{phi}$ is quite good. As the noise level increases, $\Psi_{n}^{dcor}$ takes the lead. The overall performance of $\Psi_{n}^{cor}$ is competitive except for the ‘Circle $\lambda$’ case. HHG performed well but it lagged behind our proposed tests most of the time. HSIC performed well in ‘Parabola $\lambda$’ and ‘Circle $\lambda$’ distributions, but didn’t achieve satisfactory power in other distributions. dCor had somewhat competitive power in ‘Parabola $\lambda$’, but not in other distributions. MIC had low power overall, except for higher sample sizes in the ‘Sine $\lambda$’ distributions. Xicor also performed poorly except for ‘Sine $\lambda$’, where it performed the best. This is not unexpected from xicor, as it is mainly designed to detect whether $Y$ is a function of $X$. BET performed somewhat well in ‘Laminscate $\lambda$’ distribution.
5 Performance on real dataset
To demonstrate the application of our proposed multiscale framework, we considered an interesting dataset known as Galton’s peas dataset. This dataset is available under the name peas in the R package psychTools. Collected in 1875, this dataset comprises 700 observations of the average diameters of sweet peas from both mother plants and their offspring. An intriguing aspect of this dataset, pointed out by Chatterjee (2021), is that the mean diameter of peas produced by a mother plant can be viewed as a function of the mean diameter of their offspring. However, the opposite does not hold. When tested on the full dataset, our proposed tests, as well as other competing methods, rejected the null hypothesis of independence. To compare the performance of these test methods, we evaluated their empirical power using random subsets of the dataset. Four sample sizes were considered: 10, 20, 30, and 40. For each sample size, we drew random samples without replacement (SRSWOR) from the full dataset and applied all test methods. This step was repeated independently 1,000 times to compute the empirical power of each method. As in the previous section, we fixed the number of permutations at 1,000 and set the significance level at 5%. Table 6 summarizes the empirical power of all tests across the different sample sizes.
$n$ $\Psi_{n}^{phi}$ $\Psi_{n}^{cor}$ $\Psi_{n}^{dcor}$ HHG HSIC dCor MIC XICOR BET
10 0.151 0.238 0.218 0.191 0.113 0.165 0.041 0.128 0.313
20 0.272 0.554 0.565 0.360 0.175 0.271 0.197 0.197 0.614
30 0.463 0.858 0.870 0.604 0.233 0.426 0.373 0.231 0.806
40 0.654 0.980 0.981 0.783 0.314 0.556 0.714 0.286 0.921
Table 6: The empirical power of different test methods on Galton’s pea dataset.
The results indicate that $\Psi_{n}^{cor}$ and $\Psi_{n}^{dcor}$ outperformed all other methods for sample sizes of 30 and 40. BET also performed very well. HHG ranked next in performance. Despite the presence of a huge number of ties in the dataset, $\Psi_{n}^{phi}$ demonstrated good performance. Other methods lagged behind the ones mentioned above.
6 A visual tool
Our proposed method works by collecting information on the magnitude of dependence present within the neighborhoods of different sizes around the observation points. The Z-scores $z_{1},\ldots,z_{n-1}$ calculated based on the sample observations provide insights into the strength of dependence across these different neighborhood sizes. This is already visually illustrated in Figures 5 and 6 for the distributions described in Table 1 and in Figure 4. Similarly, we can exploit this visual tool to real-world datasets for visualizing the strength of dependence across these different neighborhood sizes.
We illustrate this using the yeast gene expression dataset, initially examined by Spellman et al. (1998). The original study analyzed transcript levels of 6,223 yeast genes across 23 consecutive time points. For our analysis, we used the R package minerva, which provides a revised version of the dataset with 4,381 genes, excluding those with missing observations. For this illustration, we selected four well-behaved genes—‘YBL003C’, ‘YDR224C’, ‘YGR004C’, and ‘YHR218W’—which were also chosen by Chatterjee (2021) to demonstrate XICOR’s effectiveness in detecting functional associations. Additionally, we selected four genes—‘YBR198C’, ‘YBR300C’, ‘YEL049W’, and ‘YEL057C’—which rank among those with the lowest XICOR values. For these eight genes, we calculated the Z-scores $z_{1},\ldots,z_{22}$ based on $T^{phi}$ statistics. The expression of these genes and corresponding Z-scores are plotted in Figure 9. We applied a 4-point moving average to smooth the Z-score plot for better visualization.
Figure 9: Gene expressions for a few selected genes from yeast gene expression data (Spellman et al., 1998), along with their moving average Z-score plots.
For the four well-behaved genes, it is evident that smaller neighborhoods capture a significant amount of dependence information. However, as the neighborhood size increases, the amount of information gradually decreases. In contrast, for the other four genes, dependence information remains absent across all neighborhood sizes.
@@ 7 Discussion and conclusion
Figure 10: The scatter plots of ‘Doppler’, ‘$BEX_{2}$’, ‘Lissajous curve A’, ‘Lissajous curve B’, ‘Rose curve’, ‘Spiral’, ‘Tilted square’ and ‘Five clouds’ distributions.